Using the​ CAPM, estimate the appropriate required rate of return for the three stocks listed​ here, given that the​ risk-free rate is 6 percent and the expected return for the market is 14 percent.

STOCK

BETA

A

0.62

B

1.09

C

1.48

a. Using the​ CAPM, the required rate of return for stock A is

​%.

​(Round to two decimal​ places.)

b. Using the​ CAPM, the required rate of return for stock B is

​%.

​(Round to two decimal​ places.)

c. Using the​ CAPM, the required rate of return for stock C is

​%.

​(Round to two decimal​ places
Answer #1

a

As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 6 + 0.62 * (14 - 6)
Expected return% = 10.96

b

As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 6 + 1.09 * (14 - 6)
Expected return% = 14.72

c

As per CAPM
expected return = risk-free rate + beta * (expected return on the market - risk-free rate)
Expected return% = 6 + 1.48 * (14 - 6)
Expected return% = 17.84